OpenStudy (anonymous):

y=logx If y=10, then what is x?

Parth (parthkohli):

log base = ?

OpenStudy (anonymous):

it just say x

hartnn (hartnn):

10=log x can u convert this in exponential form?

Parth (parthkohli):

Okay, so what do you get in the equation when you plug \(\rm y = 10\)?

OpenStudy (anonymous):

yes @hartnn

OpenStudy (anonymous):

10 = log x x = 10^10 you raise both powers to whatever the base of the log is to get rid of the log

hartnn (hartnn):

what u get?

OpenStudy (anonymous):

so then it will be one?

Parth (parthkohli):

People, why don't you just help her?

Parth (parthkohli):

@latremese40 Yes!

hartnn (hartnn):

no

OpenStudy (anonymous):

no, x does not equal 1

OpenStudy (anonymous):

thank you @Parthkohli

Parth (parthkohli):

Whenever you have such a form, then I prefer the exponential form.

OpenStudy (anonymous):

@ParthKohli if x and y were reversed, this would be true

hartnn (hartnn):

its not 1

Parth (parthkohli):

@latremese40 You're welcome :)

OpenStudy (anonymous):

x = 10^10

OpenStudy (anonymous):

how @hartnn

Parth (parthkohli):

@latremese40 Let me show you.

OpenStudy (anonymous):

if you do log(1) you dont get 10

Parth (parthkohli):

\[\rm 10 = \log (x)\]Right? But here, the base is 10.

OpenStudy (anonymous):

oka Parthkohli

OpenStudy (anonymous):

so raise both sides to the 10th power to get rid of the log with base 10

Parth (parthkohli):

\[\rm 10 = \log_{10}(x)\]Do you know what the exponential form is?

hartnn (hartnn):

the exponential form \(\huge if, a=log x\implies x=10^a \)

hartnn (hartnn):

so what will be 10=log x ?

OpenStudy (anonymous):

but its y=logx then you get 10^x x=1 right?

Parth (parthkohli):

@latremese40 Let's do something clever.

hartnn (hartnn):

when y=10, u get 10=log x

Parth (parthkohli):

Let's first solve for \(\rm x\).

OpenStudy (anonymous):

@ParthKohli if you plug in 1 for x you dont get 10

hartnn (hartnn):

and u get x=10^y

OpenStudy (anonymous):

how bout @parthkohli inbx me cuz im getin confused

Parth (parthkohli):

\[\rm y = \log_{10}x \]So,\[\rm 10^y = x\]Right?

Parth (parthkohli):

Now if you have y = 10, then you have\[\rm 10^{10} =x \]Got my point?

OpenStudy (anonymous):

yes i do

Parth (parthkohli):

And that's your answer! If you prefer the numeric notation, then add ten zeroes in front of 1.

hartnn (hartnn):

sorry for the confusion.

OpenStudy (anonymous):

oka... its cool hartnn

OpenStudy (anonymous):

im still confused...how is x=1 like you said earlier?

Parth (parthkohli):

@etemplin: That answer was out of confusion, so don't mind it. :)

OpenStudy (anonymous):

wow that makes sense now. so x = 10^10 right? just making sure

OpenStudy (anonymous):